Home Work & 12 -153 1 Find the determinant of 2 Find the metro 31 Find the re c itati
Answer #1

Given – soll 6.5 27 47 st) 3.1 -1.8 9.5) A) 19.4 63.71 4.2 SH 1.1 3) sal sol - 2.415.4)(5-5) - CI-D) 6-1-8)) - (-6-5)/4-7)(5-A2 = -1-6.5 3.7 T18 s. - - ((6.5) (5.5) + (-8)(3-))) 2.4 8.4 3.71-89.09. A2 = 3.7 3./ S5 - 04)(-5) +(3-7) (3-1) - 24.67 A23 -steps: from ean & Ta = – jI b + IC Ia = - ( 70+ Hoj) + 30 +10) 70 - Hoj +30 +loj - - 10j -401-1) + 30+10j - 10-601 Ia 92.191-